OpenStudy (anonymous):

fan and medal. need help quick! show the steps plz! write the equation of the function g(x) if g(x)=f(x-3)+1 and f(x)=x^3+1.

OpenStudy (johnweldon1993):

\[\large f(x) = x^3 + 1\] \(\large f(x-3)\) means replace every 'x' you see in f(x) with (x-3) so that would give you \[\large f(x - 3) = (x-3)^3 + 1\] Now since g(x) = f(x-3) + 1 that means \[\large g(x) = f(x-3)+1 = (x-3)^3 + 1 + 1\]

OpenStudy (anonymous):

so this would be my final answer then?o.o

OpenStudy (johnweldon1993):

Well obviously we can combine the 1+1 lol \[\large g(x) = (x-3)^3 + 2\] Now as far as the "final" answer...this is technically correct...however if you wish to expand the polynomial you absolutely can :)

OpenStudy (anonymous):

oh thank you so much!!!!! :)

OpenStudy (johnweldon1993):

Not a problem :)